OpenStudy (solomonzelman):

I need help on simplifying a fraction. TRIG. I need the fastest way to do this. I am usually able to do them, but this one I got stock on.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{Csc(x)+Cot(x) }{ Csc(x)-Cot(x)} }\) the first thing I did, is that I put everything in terms of sines and cosines, add the fractions that I got on top and bottom, and when I divided them I canceled Sin(x) on top and bottom. I got (now, as my result) \(\Huge\color{blue}{ \sf \frac{1+Cos(x) }{ 1-Cos(x)} }\) and then I broke my head (dky)

OpenStudy (mathmale):

Try multiplying both numerator and denominator by (1+cos x) and see where that gets you.

OpenStudy (solomonzelman):

Yeah, I tried to do that (ik this is conjugate) \(\Huge\color{blue}{ \bf \frac{Cos^2(x)+2Cos(x)+1}{Sin^2(x)} }\)

OpenStudy (solomonzelman):

Well, or 1-Cos^2x

OpenStudy (solomonzelman):

I got disconnected.

OpenStudy (raden):

(1+cos(x))/(1-cos(x)) then, you can use these identities : 1+cosx= 2cos^2 (x/2) and 1-cosx= 2sin^2 (x/2)

OpenStudy (solomonzelman):

I got disconnected, but I think I know how to do it...

OpenStudy (mathmale):

Try leaving the numerator as (1+cos x)^2. This may or may not help. There could be multiple ways of presenting your final result. Choose the one that you think is simplest.

terenzreignz (terenzreignz):

Oh, logic, why must you be ambiguous? :D What exactly are we supposed to be doing here? How simple must simple be? Like just one trigonometric function?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf \frac{1+cos(x)}{1-cos(x)} }\) first I am going to use conjungate, (denominator is Sin^2(x) ) and multiply everything times Csc^2(x) saying that I will have nothing on the bottom \(\large\color{blue}{ \bf Csc^2(x)~~(~Cos^2(x)+2Cos(x)+1~)}\) some monipulation. \(\large\color{blue}{ \bf Csc^2(x)~~(~Cos^2(x)+2Cos(x)+Sin^2(x)+Cos^2(x)~~~)}\) \(\large\color{blue}{ \bf Cot^2(x)+Cot^2(x)+ 2Cos(x)Csc^2(x)+1}\) \(\large\color{blue}{ \bf 2Cot^2(x)+ 2Cot(x)Csc(x)+1}\) idk, it said as simple as possible

terenzreignz (terenzreignz):

We might have very different ideas on what constitute simple ;) Although, I'd rather work exclusively with sines and cosines, for reasons I'm not really sure of :D

terenzreignz (terenzreignz):

Okay, here's the deal... \[\Large \cot\left(\frac x 2\right)\] while looking particularly nasty... is going to be your best friend here :)

OpenStudy (solomonzelman):

half angle? I really haven't thought about that -:(

OpenStudy (raden):

it should be cot^2 (x/2)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf 2Cot^2(x)+ 2Cot(x)Csc(x)+1}\) using 1+cot^2(x)=csc^2(x) \(\large\color{blue}{ \bf Cot^2(x)+Csc^2(x)+ 2Cot(x)Csc(x)}\) This is what my next step would have been. yeah, cot^2 , right.

terenzreignz (terenzreignz):

\[\Large \cot \frac x 2 = \frac{\cos \frac x2}{\sin\frac x2}= \frac{\sqrt{\frac{1-\cos(x)}{2}}}{\sqrt{\frac{1+\cos(x)}{2}}}\] And d****t, @RadEn a spoiler alert next time? XD

OpenStudy (solomonzelman):

I understand you are truly the best

terenzreignz (terenzreignz):

Wait a minute... mixed up, hang on \[\LARGE\cot \frac x 2 = \frac{\cos \frac x2}{\sin\frac x2}= \frac{\sqrt{\frac{1+\cos(x)}{2}}}{\sqrt{\frac{1-\cos(x)}{2}}}\] that's better...

terenzreignz (terenzreignz):

lol Zelman, appreciate the kudos, but I didn't mean it like that? ^_^ Peace :)

OpenStudy (solomonzelman):

I got the idea, thank you!

OpenStudy (raden):

i just remember 2 identities below cos^2 (a) = (1+cos(2a)/2 and cos^2 (a) = (1+cos(2a)/2 actually, by manipulation we can take that a = x/2

OpenStudy (raden):

oppsss. mistake in vaste :) cos^2 (a) = (1+cos(2a)/2 and sin^2 (a) = (1-cos(2a)/2

OpenStudy (solomonzelman):

Yeah, I get it.... I like how terenzreignz did it. It was just awesome :)))

terenzreignz (terenzreignz):

...yeah... I suppose. :> I have a name, it's TJ (not really), so use it next time, ok, Zelman? ^^ Peace ^_^

OpenStudy (solomonzelman):

Yeah, terenzreignz :) Thank you for your help !

terenzreignz (terenzreignz):

^mother of irony XD

OpenStudy (solomonzelman):

Wait, what ?

terenzreignz (terenzreignz):

I meant don't call me terenzreignz, actually :) No real biggie, more like a pet peeve ^_^

OpenStudy (solomonzelman):

Well, you called yourself that, but whatever you want, pet peeve

terenzreignz (terenzreignz):

<sigh> XD From now on, just stick to "Terenz" I just logged in to OS to help in this question... so, signing off now :D --------- Terenz out

OpenStudy (solomonzelman):

K